# Deep Reinforcement Learning for Real-Time Drone Routing in Post-Disaster Road Assessment Without Domain Knowledge

### v1

@@ 4.1 Transformation method
This problem involves two distinct networks: the road network $G=(N,A)$ and the fully connected node network $G^{\prime}=(N,A^{\prime})$. As visualized in Figure 2, (a) depicts the original road network $G$, where nodes are connected only via physical road links (e.g., edges between adjacent intersections), forming the structure where drones conduct damage assessments. (b) shows the fully connected auxiliary network $G^{\prime}$, in which every node is directly linked to all other nodes, representing the direct transit paths drones can take to bypass road links. However, their integration into a dual-network configuration $G^{\prime\prime}=(N,A+A^{\prime})$ introduces ambiguities, as shown in (c): multiple edges (road links vs. direct links) may connect the same pair of nodes, such as edge $(1,2)$, which could represent either a road link from $G$ or a direct link from $G^{\prime}$.
To resolve this, existing studies have proposed transforming link-based routing networks into equivalent node-based ones. For example, Pearn (1988) converted a network with $|A|$ links into an equivalent structure with $3|A|+1$ nodes, while Baldacci and Maniezzo (2006) and Longo et al. (2006) developed methods to reduce this to $2|A|+1$ nodes. However, these approaches often generate redundant links, increasing computational complexity. For the drone path planning problem, we adopt a traffic assignment method using intersection construction (Boyles et al. 2023) to reformulate the problem as a node-based routing task. As shown in Figure 3 (a), each road link $(i,j)\in A$ is split by an artificial node $p\in\mathcal{P}$, creating two new links: $(i,p)$ and $(p,j)$. This transforms the dual network $G^{\prime\prime}=(N,A+A^{\prime})$ into a reduced network $\bar{G}=(\bar{N},\bar{A})=(N\cup\mathcal{P},2A\cup A^{\prime})$, where $|\mathcal{P}|=|A|$ (see Figure 3 (b)). In $\bar{G}$, the ambiguity of multiple edges between nodes is resolved: traversing $(i,p)$ and $(p,j)$ corresponds to assessing the original link $(i,j)$, while direct links in $A^{\prime}$ remain unchanged. By setting the information value of $p$ as $c_{p}=c_{ij}$ and splitting the traversal time as $t_{ip}=t_{pj}=t_{ij}/2$, the link-based routing problem is converted into a node-based one (examples in Figure 3 (c) and Figure 3 (d)). Notably, since $|A|\ll|A^{\prime}|+|N|$ (Boyles et al. 2023), this transformation adds only $|A|$ nodes and $|A|$ links, resulting in lower complexity compared to existing methods (Pearn 1988, Baldacci and Maniezzo 2006, Longo et al. 2006).
Figure 2: Dual Network Ambiguity: Overlapping Links Between Nodes
Figure 3: Link-to-Node Transformation via Artificial Nodes (where 5,6,7,and 8 are aritifical nodes.)
@@ 4.2 Coordinate generation
To systematically solve the problem, we address the coordinate generation for artificial nodes. In a given road network, geographical coordinates of original nodes are predefined (as shown in Figure 3 (c)), along with the corresponding road lengths, but coordinates of artificial nodes (e.g., nodes 5, 6, 7, 8 in Figure 3 (d)) are not. We propose a method to determine their coordinates as follows: Consider two nodes $A(X_{A},Y_{A})$ and $B(X_{B},Y_{B})$ in a 2D plane, with link $(A,B)$ of length $L$. We aim to find coordinates of point $C(X_{C},Y_{C})$ such that $AC=BC=L/2$ and $C$ lies on the perpendicular bisector of $AB$ (see Figure 4). Compute the midpoint $D$ of $AB$: $X_{D}=(X_{A}+X_{B})/2$, $Y_{D}=(Y_{A}+Y_{B})/2$. The Euclidean distance between $A$ and $B$ is $d_{AB}=\sqrt{(X_{B}-X_{A})^{2}+(Y_{B}-Y_{A})^{2}}$. Since $C$ lies on the perpendicular bisector of $AB$, the direction vector of this bisector is $(-(Y_{B}-Y_{A}),X_{B}-X_{A})$, and its unit normal vector is $(-(Y_{B}-Y_{A})/d_{AB},(X_{B}-X_{A})/d_{AB})$. The coordinates of $C$ are derived as (detailed derivation in Appendix A. Derivation of point $C$ coordinates):
@@
Figure 4: Artificial Node Coordinate Calculation: Perpendicular Bisector Method
A key insight is that, via the proposed transformation (converting link-based routing to node-based routing), this drone routing problem resembles the orienteering problem (OP)—where nodes have prizes, and the goal is to maximize total collected prizes while keeping route length within a maximum limit (Gunawan et al. 2016, Kool et al. 2018, Zhang et al. 2023). However, critical differences exist as listed below, precluding direct application of existing OP algorithms. First, original nodes in $N$ carry no information value, while artificial nodes in $\mathcal{P}$ (representing road links) do. Second, in the transformed network $\bar{G}$, all nodes in $N$ are mutually connected, but artificial nodes in $\mathcal{P}$ are isolated from each other—each connects only to two specific nodes in $N$. Third, and most notably, artificial nodes in $\mathcal{P}$ are exclusive: each can be visited by only one drone (as a road link’s information value is collected once), whereas nodes in $N$ (valueless) can be revisited. For example, in Figure 3 (b) with four drones (each limited to one link by time constraints), if nodes 2 and 3 were non-revisitable, only two links’ information could be collected—suboptimal compared to four with revisits allowed. Additionally, revisits enable loops in drone paths: in Figure 5, a single drone’s optimal route (1,5,2,3,7,2,6,4) includes a loop (2,3,7,2), which would be impossible under non-revisitable constraints. This contrasts with OP’s Miller-Tucker-Zemlin (MTZ) constraints (Miller et al. 1960), which prohibit cycles to ensure path continuity. These structural differences mean existing OP algorithms (exact or heuristic) cannot be directly applied. This highlights a common limitation of traditional algorithmic solutions: even minor modifications to the problem structure often require substantial redesign, which can be time-consuming and heavily reliant on expert domain knowledge. Moreover, in disaster-response scenarios where real-time decision-making is critical, there is an urgent need for methods that can deliver high-quality solutions as swiftly as possible, ideally in real time. In the following sections, we present a novel approach designed to tackle these challenges.
@@ 5 Attention-based encoder-decoder model
To address the real-time solution requirement and eliminate the need for domain knowledge in algorithm design, we propose a learning-based framework, where a deep neural network is employed to map problem instance input data (denoted as vector $\mathbf{x}$) to a valid solution (denoted as vector $\mathbf{y}$). Given the absence of large-scale labeled training data $(\mathbf{x},\mathbf{y})$ for this novel problem—a common challenge for many emerging combinatorial optimization problems—we adopt reinforcement learning (RL) to train the deep neural network, enabling real-time inference without domain-specific expertise.
@@ 5.1 Encoder
Leveraging the transformer architecture (Vaswani et al. 2017), the encoder employs embedding techniques to process and represent input data features. In the initial encoding stage, an embedding layer maps raw node features into a higher-dimensional latent space. Nodes are partitioned into two distinct categories:
@@
This heterogeneous embedding strategy differentiates node types and embeds global constraints into the depot’s representation, forming a holistic input for subsequent attention-based processing. These embedding are then processed by $l$ transformer layers (Kool et al. 2018), each consisting of two sublayers: a multi-head attention (MHA) sublayer and a node-wise fully connected feed-forward network (FFN) sublayer. Inspired by residual networks, a residual connection is employed to mitigate the issues of vanishing or exploding gradients in deep neural networks (He et al. 2016). Additionally, instance normalization (IN) is applied to accelerate training and enhance model performance. Mathematically, a transformer layer is defined as:
@@ 5.2 Decoder
where $C$ serves as a clipping parameter for the tanh function, enhancing solution exploration in the search space (Bello et al. 2016, Berto et al. 2024). Upon obtaining the probability distribution $p_{i}$, the current time step $t$ is finalized, and the state is updated to $t+1$, allowing the decoder to construct solutions autoregressively. The decoder terminates when either of the following conditions is met, yielding a feasible solution:
@@
Given that the detailed architectures of MHA, IN, FFN, and SHA represent mature techniques in large language models—thoroughly discussed in existing literature such as Vaswani et al. (2017), Kool et al. (2018), Kwon et al. (2020), Luo et al. (2023), Zhou et al. (2024), Liu et al. (2024a), Berto et al. (2024)—this study refrains from redundant technical elaboration.
@@ 5.3 Training method
The stochastic policy in equation (4) is parameterized by $\theta$ (encompassing all trainable variables of AEDM). For model optimization, policy gradient methods iteratively refine $\theta$ via gradient estimation of the expected reward. While Kool et al. (2018) proposed the REINFORCE estimator—using a deterministic greedy rollout of the current best policy as baseline (updated periodically via paired t-test with $\alpha=5\%$)—recent work by Kwon et al. (2020) showed that policy optimization with multiple optima (POMO) outperforms REINFORCE. Thus, we train AEDM using POMO.
Notably, our problem diverges from the original POMO in four key aspects. First, fixing $K$ (drone number), $p_{\max}$ (assessment time limit), and $Q$ (battery constraint) reduces the problem to single-task learning—analogous to training models for the fixed-vehicle number vehicle routing problem in Zhang et al. (2020). While simplifying training, this approach necessitates separate models for each parameter combination and requires matching models to combinations during deployment, which is inefficient. Inspired by multi-task capabilities of LLM, we advance AEDM as a multi-task framework: it is trained to handle diverse $(K,p_{\max},Q)$ combinations and, crucially, generalizes to unseen ones.
Second, two strategies exist for training such a multi-task model: (1) Directly mixing diverse parameter combinations within individual batches (Berto et al. 2024); (2) Training a fixed parameter set per batch, with varying combinations across successive batches (Zhou et al. 2024, Liu et al. 2024a). For our problem, decoder output lengths vary significantly across different combinations of $K$, $p_{\max}$, and $Q$. For instance, with fixed $p_{\max}$ and $Q$, the decoder output length for $K=4$ is nearly double that for $K=2$. The latter accelerates training by ensuring near-uniform decoder output lengths across instances within each batch. Thus, we adopt the second strategy.
@@
where $\mu_{t}^{(pc)}$ is the mean reward of the current batch for $pc$, $\hat{\mu}_{t-1}^{(pc)}$ is the smoothed mean from the previous batch, and $\gamma\in[0,1]$ (set to 0.25 following Berto et al. (2024)) controls the smoothing intensity, with initialization $\hat{\mu}_{1}^{(pc)}=\mu_{1}^{(pc)}$. We then calculate the exponentially smoothed variance to capture reward dispersion:
@@
where $\epsilon\approx 10^{-8}$ avoids division by zero. This approach simultaneously tracks changes in both mean and variance across batches, ensuring stable multi-task convergence by anchoring rewards to their historical trajectories within each parameter regime. As validated by the ablation study in Appendix C. Ablation study, this normalization strategy outperforms alternatives such as batch-wise Z-score normalization and no normalization, minimizing optimality gaps and enhancing model stability across diverse parameter combinations.
Fourth, for POMO implementation (Kwon et al. 2020), we sample a solution set $\Pi$, where each $\pi\in\Pi$ is determined by AEDM. Given the drone routing problem’s nature—all drones depart from a shared depot—we enhance sample diversity by sequentially selecting $N$ nodes from the transformed road network as a secondary visit set. Due to the network’s structure: nodes with non-zero damage information values are disconnected from the depot, while those with zero values are connected. Thus, $N$ nodes are sequentially chosen to form the secondary visit set for drones.
@@
Here, $s$ denotes a specific problem instance; $R(\pi|s)$ represents the reward yielded by solution $\pi$, corresponding to the total collected damage information value. The shared baseline $b(s)$ is defined as $b(s)=\frac{1}{|\Pi|}\sum_{\pi\in\Pi}R(\pi|s)$. Finally, $p_{\theta}(\pi|s)$ denotes the aggregated selection probabilities of nodes at each decoder step. A summary of the AEDM training procedure is provided in Algorithm 1. Additionally, the model’s learning mechanism—transitioning from memorizing input parameters to reasoning about problem structures—is further validated by the t-distributed stochastic neighbor embedding (t-SNE) interpretability analysis in Appendix D. Method interpretability, which illustrates the evolutionary trajectory of features across encoder layers and confirms the effectiveness of the proposed training framework.
@@ 5.4.1 Training instance generation
In contrast to prior DRL studies focused on combinatorial optimization problems such as vehicle routing problem (VRP) and traveling salesman problem (TSP) (Kool et al. 2018, Kwon et al. 2020, Luo et al. 2023, Liu et al. 2024a, Berto et al. 2024)—which generate training instances via random node placement in a coordinate system—our drone routing problem requires road network-based instances. Given the scarcity of public road network datasets (fewer than 20, e.g., https://github.com/bstabler/TransportationNetworks), insufficient for effective DRL training, we propose a synthetic road network generation method.
@@
    Link Pruning: Randomly remove edge subsets (prioritizing boundary nodes) while maintaining connectivity, resulting in $A$ edges (see Figure 7 (b)).
@@
This method generates large-scale, diverse, and realistic road network instances—addressing dataset scarcity and enabling effective DRL training for disaster-response drone routing.
Figure 7: The three-stage synthetic road network generation for drone routing training. (a) Initializes a $N=30$ node grid in $0,1]^{2}$. (b) Prunes edges (prioritizing boundaries) to retain $A=36$ edges while ensuring connectivity, mimicking real-world sparsity. (c) Applies bounded random node perturbations within $0,1]^{2}$ to break regularity and introduce topological diversity. This pipeline addresses dataset scarcity, enabling realistic training environments for DRL-based drone routing optimization.
5.4.2 Testing intance generation
Given the transformed road network structure, we select $N$ nodes as the secondary visit set for drones. A critical limitation of POMO’s multi-greedy inference for our problem is that $N$ (the number of greedy rollouts) cannot scale arbitrarily. Inspired by data augmentation in computer vision (e.g., cropping and flipping) and natural language processing (e.g., back-translation), Kwon et al. (2020) proposed generating new instances via coordinate transformations.

### v2

@@ 4.1 Transformation method
This problem involves two distinct networks: the road network $G=(N,A)$ and the fully connected node network $G^{\prime}=(N,A^{\prime})$. As visualized in Figure 2, (a) depicts the original road network $G$, where nodes are connected only via physical road links (e.g., links between adjacent intersections), forming the structure where drones conduct damage assessments. (b) shows the fully connected auxiliary network $G^{\prime}$, in which every node is directly linked to all other nodes, representing the direct transit paths drones can take to bypass road links. However, their integration into a dual-network configuration $G^{\prime\prime}=(N,A+A^{\prime})$ introduces ambiguities, as shown in (c): multiple links (road links vs. direct links) may connect the same pair of nodes, such as link $(1,2)$, which could represent either a road link from $G$ or a direct link from $G^{\prime}$.
To resolve this, existing studies have proposed transforming link-based routing networks into equivalent node-based ones. For example, pearn1988new converted a network with $|A|$ links into an equivalent structure with $3|A|+1$ nodes, while baldacci2006exact and longo2006solving developed methods to reduce this to $2|A|+1$ nodes. However, these approaches often generate redundant links, increasing computational complexity. For the drone path planning problem, we adopt a traffic assignment method using intersection construction (blubook) to reformulate the problem as a node-based routing task. As shown in Figure 3 (a), each road link $(i,j)\in A$ is split by an artificial node $p\in\mathcal{P}$, creating two new links: $(i,p)$ and $(p,j)$. This transforms the dual network $G^{\prime\prime}=(N,A+A^{\prime})$ into a reduced network $\bar{G}=(\bar{N},\bar{A})=(N\cup\mathcal{P},2A\cup A^{\prime})$, where $|\mathcal{P}|=|A|$ (see Figure 3 (b)). In $\bar{G}$, the ambiguity of multiple links between nodes is resolved: traversing $(i,p)$ and $(p,j)$ corresponds to assessing the original link $(i,j)$, while direct links in $A^{\prime}$ remain unchanged. By setting the information value of $p$ as $c_{p}=c_{ij}$ and splitting the traversal time as $t_{ip}=t_{pj}=t_{ij}/2$, the link-based routing problem is converted into a node-based one (examples in Figure 3 (c) and Figure 3 (d)). Notably, since $|A|\ll|A^{\prime}|+|N|$ (blubook), this transformation adds only $|A|$ nodes and $|A|$ links, resulting in lower complexity compared to existing methods (pearn1988new; baldacci2006exact; longo2006solving).
Figure 2: Dual Network Ambiguity with Overlapping Links Between Nodes
Figure 3: Link-to-Node Transformation via Artificial Nodes (where 5,6,7,and 8 are artificial nodes.)
@@ 4.2 Coordinate generation
To systematically solve the problem, we address the coordinate generation for artificial nodes. In a given road network, geographical coordinates of original nodes are predefined (as shown in Figure 3 (c)), along with the corresponding road lengths, but coordinates of artificial nodes (e.g., nodes 5, 6, 7, 8 in Figure 3 (d)) are not. We propose a method to determine their coordinates as follows: Consider two nodes $A(X_{A},Y_{A})$ and $B(X_{B},Y_{B})$ in a 2D plane, with link $(A,B)$ of length $L$. We aim to find coordinates of point $C(X_{C},Y_{C})$ such that $AC=BC=L/2$ and $C$ lies on the perpendicular bisector of $AB$ (see Figure 4). Compute the midpoint $D$ of $AB$: $X_{D}=(X_{A}+X_{B})/2$, $Y_{D}=(Y_{A}+Y_{B})/2$. The Euclidean distance between $A$ and $B$ is $d_{AB}=\sqrt{(X_{B}-X_{A})^{2}+(Y_{B}-Y_{A})^{2}}$. Since $C$ lies on the perpendicular bisector of $AB$, the direction vector of this bisector is $(-(Y_{B}-Y_{A}),X_{B}-X_{A})$, and its unit normal vector is $(-(Y_{B}-Y_{A})/d_{AB},(X_{B}-X_{A})/d_{AB})$. The coordinates of $C$ are derived as (detailed derivation in Appendix A):
@@
Figure 4: Artificial Node Coordinate Calculation Using Perpendicular Bisector Method
A key insight is that, via the proposed transformation (converting link-based routing to node-based routing), this drone routing problem resembles the orienteering problem (OP), where nodes have prizes, and the goal is to maximize total collected prizes while keeping route length within a maximum limit (gunawan2016orienteering; kool2018attention; zhang2023robust). However, critical differences exist as listed below, precluding direct application of existing OP algorithms. First, original nodes in $N$ carry no information value, while artificial nodes in $\mathcal{P}$ (representing road links) do. Second, in the transformed network $\bar{G}$, all nodes in $N$ are mutually connected, but artificial nodes in $\mathcal{P}$ are isolated from each other; each connects only to two specific nodes in $N$. Third, and most notably, artificial nodes in $\mathcal{P}$ are exclusive: each can be visited by only one drone (as a road link’s information value is collected once), whereas nodes in $N$ (valueless) can be revisited. For example, in Figure 3 (b) with four drones (each limited to one link by time constraints), if nodes 2 and 3 were non-revisitable, only two links’ information could be collected, which is suboptimal compared to four with revisits allowed. Additionally, revisits enable loops in drone paths: in Figure 5, a single drone’s optimal route (1,5,2,3,7,2,6,4) includes a loop (2,3,7,2), which would be impossible under non-revisitable constraints. This contrasts with OP’s Miller-Tucker-Zemlin (MTZ) constraints (miller1960integer), which prohibit cycles to ensure path continuity. These structural differences mean existing OP algorithms (exact or heuristic) cannot be directly applied. This highlights a common limitation of traditional algorithmic solutions: even minor modifications to the problem structure often require substantial redesign, which can be time-consuming and heavily reliant on expert domain knowledge. Moreover, in disaster-response scenarios where rapid decision-making is critical, there is an urgent need for methods that can deliver high-quality solutions as swiftly as possible, ideally in real time. In the following sections, we present a novel approach designed to tackle these challenges.
@@ 5 Attention-based encoder-decoder model
To address the rapid solution requirement and eliminate the need for domain knowledge in algorithm design, we propose a learning-based framework, where a deep neural network is employed to map problem instance input data (denoted as vector $\mathbf{x}$) to a valid solution (denoted as vector $\mathbf{y}$). Given the absence of large-scale labeled training data $(\mathbf{x},\mathbf{y})$ for this novel problem, which is a common challenge for many emerging combinatorial optimization problems, we adopt reinforcement learning (RL) to train the deep neural network, enabling rapid inference without domain-specific expertise.
@@ 5.1 Encoder
Leveraging the transformer architecture (waswani2017attention), the encoder employs embedding techniques to process and represent input data features. In the initial encoding stage, an embedding layer maps raw node features into a higher-dimensional latent space. Nodes are partitioned into two distinct categories:
@@
This heterogeneous embedding strategy differentiates node types and embeds global constraints into the depot’s representation, forming a holistic input for subsequent attention-based processing. These embedding are then processed by $l$ transformer layers (kool2018attention), each consisting of two sublayers: a multi-head attention (MHA) sublayer and a node-wise fully connected feed-forward network (FFN) sublayer. Inspired by residual networks, a residual connection is employed to mitigate the issues of vanishing or exploding gradients in deep neural networks (he2016identity). Additionally, instance normalization (IN) is applied to accelerate training and enhance model performance. Mathematically, a transformer layer is defined as:
@@ 5.2 Decoder
where $C$ serves as a clipping parameter for the tanh function, enhancing solution exploration in the search space (bello2016neural; berto2024routefinder). Upon obtaining the probability distribution $p_{i}$, the current time step $t$ is finalized, and the state is updated to $t+1$, allowing the decoder to construct solutions autoregressively. The decoder terminates when either of the following conditions is met, yielding a feasible solution:
@@
Given that the detailed architectures of MHA, IN, FFN, and SHA represent mature techniques in large language models, which are thoroughly discussed in existing literature such as waswani2017attention; kool2018attention; kwon2020pomo; luo2023neural; zhou2024mvmoe; liu2024multi; berto2024routefinder, this study refrains from redundant technical elaboration.
@@ 5.3 Training method
The stochastic policy in equation (4) is parameterized by $\theta$ (encompassing all trainable variables of AEDM). For model optimization, policy gradient methods iteratively refine $\theta$ via gradient estimation of the expected reward. While kool2018attention proposed the REINFORCE estimator, which uses a deterministic greedy rollout of the current best policy as baseline (updated periodically via paired t-test with $\alpha=5\%$), recent work by kwon2020pomo showed that policy optimization with multiple optima (POMO) outperforms REINFORCE. Thus, we train AEDM using POMO.
Notably, our problem diverges from the original POMO in four key aspects. First, fixing $K$ (drone number), $p_{\max}$ (assessment time limit), and $Q$ (battery constraint) reduces the problem to single-task learning, which is analogous to training models for the fixed-vehicle number vehicle routing problem in zhang2020multi. While simplifying training, this approach necessitates separate models for each parameter combination and requires matching models to combinations during deployment, which is inefficient. Inspired by multi-task capabilities of LLM, we advance AEDM as a multi-task framework: it is trained to handle diverse $(K,p_{\max},Q)$ combinations and, crucially, generalizes to unseen ones.
Second, two strategies exist for training such a multi-task model: (1) Directly mixing diverse parameter combinations within individual batches (berto2024routefinder); (2) Training a fixed parameter set per batch, with varying combinations across successive batches (zhou2024mvmoe; liu2024multi). For our problem, decoder output lengths vary significantly across different combinations of $K$, $p_{\max}$, and $Q$. For instance, with fixed $p_{\max}$ and $Q$, the decoder output length for $K=4$ is nearly double that for $K=2$. The latter accelerates training by ensuring near-uniform decoder output lengths across instances within each batch. Thus, we adopt the second strategy.
@@
where $\mu_{t}^{(pc)}$ is the mean reward of the current batch for $pc$, $\hat{\mu}_{t-1}^{(pc)}$ is the smoothed mean from the previous batch, and $\gamma\in[0,1]$ (set to 0.25 following berto2024routefinder) controls the smoothing intensity, with initialization $\hat{\mu}_{1}^{(pc)}=\mu_{1}^{(pc)}$. We then calculate the exponentially smoothed variance to capture reward dispersion:
@@
where $\epsilon\approx 10^{-8}$ avoids division by zero. This approach simultaneously tracks changes in both mean and variance across batches, ensuring stable multi-task convergence by anchoring rewards to their historical trajectories within each parameter regime. As validated by the ablation study in Appendix D, this normalization strategy outperforms alternatives such as batch-wise Z-score normalization and no normalization, minimizing relative gaps and enhancing model stability across diverse parameter combinations.
Fourth, for POMO implementation (kwon2020pomo), we sample a solution set $\Pi$, where each $\pi\in\Pi$ is determined by AEDM. Given the drone routing problem’s nature, where all drones depart from a shared depot, we enhance sample diversity by sequentially selecting $N$ nodes from the transformed road network as a secondary visit set. Due to the network’s structure, nodes with non-zero damage information values are disconnected from the depot, while those with zero values are connected. Thus, $N$ nodes are sequentially chosen to form the secondary visit set for drones.
@@
Here, $s$ denotes a specific problem instance; $R(\pi|s)$ represents the reward yielded by solution $\pi$, corresponding to the total collected damage information value. The shared baseline $b(s)$ is defined as $b(s)=\frac{1}{|\Pi|}\sum_{\pi\in\Pi}R(\pi|s)$. Finally, $p_{\theta}(\pi|s)$ denotes the aggregated selection probabilities of nodes at each decoder step. A summary of the AEDM training procedure is provided in Algorithm 1. Additionally, the model’s learning mechanism, which transitions from memorizing input parameters to reasoning about problem structures, is further validated by the t-distributed stochastic neighbor embedding (t-SNE) interpretability analysis in Appendix E, which illustrates the evolutionary trajectory of features across encoder layers and confirms the effectiveness of the proposed training framework.
@@ 5.4.1 Training instance generation
In contrast to prior DRL studies focused on combinatorial optimization problems such as vehicle routing problem (VRP) and traveling salesman problem (TSP) (kool2018attention; kwon2020pomo; luo2023neural; liu2024multi; berto2024routefinder), which generate training instances via random node placement in a coordinate system, our drone routing problem requires road network-based instances. Given the scarcity of public road network datasets (fewer than 20, e.g., https://github.com/bstabler/TransportationNetworks), which is insufficient for effective DRL training, we propose a synthetic road network generation method.
@@
    Link Pruning: Randomly remove link subsets (prioritizing boundary nodes) while maintaining connectivity, resulting in $A$ links (see Figure 7 (b)).
@@
This method generates large-scale, diverse, and realistic road network instances, addressing dataset scarcity and enabling effective DRL training for disaster-response drone routing.
Figure 7: The three-stage synthetic road network generation for drone routing training. (a) Initializes a $N=30$ node grid in $0,1]^{2}$. (b) Prunes links (prioritizing boundaries) to retain $A=36$ links while ensuring connectivity, mimicking real-world sparsity. (c) Applies bounded random node perturbations within $0,1]^{2}$ to break regularity and introduce topological diversity. This pipeline addresses dataset scarcity, enabling realistic training environments for DRL-based drone routing optimization.
5.4.2 Testing instance generation
Given the transformed road network structure, we select $N$ nodes as the secondary visit set for drones. A critical limitation of POMO’s multi-greedy inference for our problem is that $N$ (the number of greedy rollouts) cannot scale arbitrarily. Inspired by data augmentation in computer vision (e.g., cropping and flipping) and natural language processing (e.g., back-translation), kwon2020pomo proposed generating new instances via coordinate transformations.